OpenStudy (anonymous):

A constant force applied to A causes A to accelerate at 5 m/s^2 . The same force applied to B causes an acceleration of 4 m/s^2. What is the ratio mA/mB of the mass of A to the mass of B?

OpenStudy (ash2326):

revo you know the relation between force, mass and acceleration?

OpenStudy (anonymous):

yes, F=m*a

OpenStudy (ash2326):

For the first body, let mass be m1 and write the relation

OpenStudy (anonymous):

A = m1*a = m1(5 m/s^2) B = m2*b = m2(4m/s^2), right?

OpenStudy (ash2326):

yeah , correct, but you know the force applied to them is same, so assume A=B

OpenStudy (anonymous):

m1(5 m/s^2) = m2 (4 m/s^2)

OpenStudy (ash2326):

now find m1/m2, that's your answer

OpenStudy (anonymous):

is it (5 m/s^2) / (4 m/s^2) = 1.25?

OpenStudy (ash2326):

oh, you need to find m1/m2 we have m1*5=m2*4 so m1/m2=4/5=0.8

OpenStudy (anonymous):

oh, I see, thank you very much!

OpenStudy (ash2326):

welcome but you did everything, great work:)